OpenStudy (madeleine2):

A car has an initial value of $20,000. The value then "depreciates" or decreases by 10% of its initial value each year. Which equation below gives the value of the car, v, after y years? a. v = - 0.10y + 2,000 b. v = 0.10y - 2,000 c. v = 2,000y - 20,000 d. v = -2,000y + 20,000

OpenStudy (mathstudent55):

The car's original value is y. After 1 year, it lost 10% of its value. How much value did it lose in 1 year?

OpenStudy (madeleine2):

is the answer b?

OpenStudy (mathstudent55):

No.

OpenStudy (mathstudent55):

If the car is worth y, and it loses 10% of its value in 1 year, how mucgh value in terms of y does it lose in 1 year?

OpenStudy (madeleine2):

is it neither a or b becuase y is supposed to be 20,000 or is it the other way around

OpenStudy (madeleine2):

wait is the answer then c

OpenStudy (mathstudent55):

Are you going to guess the answers one by one until you get tot he correct one? I think it's better to understand it and figure out the correct answer without guesses.

OpenStudy (mathstudent55):

The car was worth $20,000 to start with. It loses in value 10% of $20,000 each year. How much does it lose in value each year? In other words, what is 10% of $20,000?

OpenStudy (madeleine2):

no i think its c becuase its subracting which means its decreasing

OpenStudy (mathstudent55):

I see a negative sign in both c and d. How can you tell it's c and not d? Please answer my question above. We'll get to the correct answer soon.

OpenStudy (mathstudent55):

Can you answer this? The car was worth $20,000 to start with. It loses in value 10% of $20,000 each year. How much does it lose in value each year? In other words, what is 10% of $20,000?

OpenStudy (madeleine2):

10% of $20,000 is 2000

OpenStudy (mathstudent55):

Correct. The car started with a value of $20,000. Each year that goes by, it loses another $2,000 in value. Ok so far?

OpenStudy (mathstudent55):

Year 0 value: 20,000 Year 1 value: 20,000 - 2000 Year 2 value: 20,000 - 2000 - 2000 = 20,000 - 2000(2) Year 3 value: 20,000 - 2000 - 2000 - 2000 = 20,000 - 2000(3) Year 4 value: 20,000 - 2000 - 2000 - 2000 - 2000 = 20,000 - 2000(4) etc. Year y value: 20,000 - 2000y

OpenStudy (mathstudent55):

Each year it loses $2000. In the first year it loses $2000 in value. In the second year it has lost 2-years' worth of $2000 a year loss, so it already lost 2 * $2000 = $4000. In the third year, it has had 3 years of losing $2000 per year, so at eh end of the 3rd year, it has lost 3 * $2000 in value. For each year that passes, multiply the number of years by $2000, and subtract from the original value of $20,000. In year y, it has lost 2000y in value. Since it started at a value of 20,000, in year y it is worth 20,000 - 2000y. 20,000 - 2000y is the same as -2000y + 20,000